# A Decade of Wheat Mapping for Lebanon

### v2

@@ I Introduction
Figure 1 shows that the proposed pipeline overcomes the stated challenges, allowing more reliable agricultural monitoring. An interactive version of this Figure can be accessed here: https://shorturl.at/PONJH.
@@ V Discussion and Future Work
TABLE II: Change in Wheat Area (km²) and Number of Wheat Fields in Lebanon Over the Last Decade
Year Wheat Area Number of Fields
2016 177.0 12,419
2017 176.8 10,604
2018 177.6 12,811
2019 254.4 17,681
2020 144.4 9,816
2021 192.7 13,602
2022 187.9 12,185
2023 224.7 14,742
2024 139.1 9,291
@@
Figure 5-(c) shows that the highest abandonment rate was in 2019-2020 ($207.7km^{2}$), closely correlated with the significant decline in the total wheat area in 2020, possibly due to economic distress. The lowest abandonment was $101km^{2}$ (2020-2021), suggesting a partial recovery or stabilization after significant disruptions.

### v3

@@ I Introduction
Figure 1 shows that the proposed pipeline overcomes the stated challenges, allowing more reliable agricultural monitoring. An interactive version of this Figure can be accessed here: https://shorturl.at/PONJH
@@ V Discussion and Future Work
TABLE II: Wheat Field Area (km²), Fields Count, and Vast Wheat Area (km²) in Lebanon, 2016–2024
Year Wheat Fields Surface Number of Fields Vast Wheat Areas
2016 247.6 15,691 75.6
2017 252.4 12,327 136.8
2018 241.9 15,635 157.4
2019 365.6 20,914 262.1
2020 203.0 12,113 50.9
2021 283.8 16,805 116.4
2022 263.1 15,049 78.9
2023 315.2 17,867 121.5
2024 199.1 11,638 63.7
Conceptually, Vast Wheat Areas quantify wheat outside parcel-like fields (e.g., diffuse plantings) thus beyond FTW’s “field” ontology; elevated values mostly signal omission.
@@
Figure 5-(c) shows that the highest abandonment rate was in 2019-2020 ($505.3km^{2}$), closely correlated with the significant decline in the total wheat area in 2020, possibly due to economic distress. The lowest abandonment was $144.7km^{2}$ (2020-2021), suggesting a partial recovery or stabilization after significant disruptions.